I don't understand this question. 9 is the difference of a number n and 7. I need to write the word sentence as an equation. And then solve the equation. I think the answer is 9-n=7. n=2?
Answer by KMST(5328)

The difference of a number n and 7 is n-7 (at least I've seen interpreted that way in books).
Since that difference is 9, n-7=9 and n=9+7, so n=16.